OpenStudy (anonymous):

how do i solve 3x + 2 < 14

OpenStudy (anonymous):

bring the two over so it 14 -2 which is 12 the divide by 3 and x is 4

OpenStudy (anonymous):

3x+2<14 3x=14-2 3x=12 x=12/3 x=<4

OpenStudy (anonymous):

MY MEDAL

OpenStudy (anonymous):

i answered first

OpenStudy (anonymous):

I have one more similar but subtraction 5x - 6 >100

OpenStudy (anonymous):

X>106/5 MY MEDAL

OpenStudy (anonymous):

Please show me how u got x>106/5 DAVIDFF i have to show my work

OpenStudy (anonymous):

just do the same again so 100 + 6 then divide it by 5

OpenStudy (anonymous):

NO MEDAL NO HELP BYE

OpenStudy (anonymous):

Samnatha so x > 20 or -20

OpenStudy (anonymous):

well on my calculator its 21.2 but if you want to round it off to 21 then its that

OpenStudy (anonymous):

Thanks Samnatha yur Awesome nd fast to

OpenStudy (anonymous):

IS 21.2

OpenStudy (anonymous):

haha no problem

OpenStudy (anonymous):

Thanks DAVIDFF yur a lifesaver to

OpenStudy (anonymous):

NO MORE HELP YOU FOR MY MEDAL BYE

OpenStudy (anonymous):

wow davidff talk about selfish

OpenStudy (anonymous):

YES MEN I HELP 2 BUT NO MY MEDAL BYE

OpenStudy (anonymous):

YES MEN I HELP 2 PEOPLE BUT NO MY MEDAL

OpenStudy (anonymous):

I am truly sorry 45 nd Goodbye